OpenStudy (anonymous):

A conducting bar moves along frictionless conducting rails connected to a 4ohm resistor. The length of the bar is 1.5m and a uniform magnetic field of 2T is applied perpendicular to the paper. What is the applied force required to move the bar to the right with a constant speed of 5m/s?

OpenStudy (anonymous):

 do u get the picture

OpenStudy (anonymous):

it seems irrelevant, but yes/

OpenStudy (aravindg):

Interesting. Obviously, emf generated is Blv. We need to relate it to force any idea?

OpenStudy (anonymous):

F=Length*I*B?

OpenStudy (aravindg):

Looks good. From the induced emf I=E/R

OpenStudy (anonymous):

Area covered by the conductor in unit time = V * L ? So Rate of change in magnetic flux = E = BLV ? You can find E when u hve value of B, L and V You are given a resistor, so probably you might want to find current:/ And you know force on a conductor perpendicular to a magnetic field = BIL Since it is at constant speed force would be equal to BIL

OpenStudy (anonymous):

F=Length*V/R*B?

OpenStudy (anonymous):

hmmm?

OpenStudy (anonymous):

How did u derive it?

OpenStudy (aravindg):

E=Blv I=BLv/R F=BIL F=B^2L^2V/R Does that look right @zaphod

OpenStudy (anonymous):

\[F = (BL)^{2} \times \frac{V}{R}\] Ofcourse

OpenStudy (aravindg):

So there you go! @Grazes :)

OpenStudy (anonymous):

yup.

OpenStudy (aravindg):

If you have more questions post them. Even I am revising magnetics atm.

OpenStudy (anonymous):

@AravindG i ve got one? and i want to discuss it wid anyone?

OpenStudy (aravindg):

Post it :) We could have a combine session for 3 of us. Will be very helpful.

OpenStudy (anonymous):

The wire in the figure carries a current I that is increasing with time at a constant rate. The induced emf in each of the loops is such that... other question: if the current is decreasing with time at a constant rate